OpenStudy (anonymous):

Suppose that r varies directly as p and inversely as q2, and that r = 27 when p = 3 and q = 2. Find r when p = 2 and q = 3.

OpenStudy (johnweldon1993):

\[\large r = \frac{kp}{q}\] r = 27 when p = 3 and q = 2 so we can solve for 'k' (our constant of variation) \[\large 27 = \frac{3k}{2}\] solve for 'k' there

OpenStudy (anonymous):

thank you

OpenStudy (johnweldon1993):

That isn't everything... Once you solve for 'k' there you will have the same equation \[\large r = \frac{kp}{q}\] And you will input p = 2, q = 3 and the 'k' you fund...and solve for 'r'

OpenStudy (anonymous):

i cant seem to find k